# On Doubly Robust Estimation with Nonignorable Missing Data Using Instrumental Variables

### v2

@@ 5 Illustration
To illustrate the proposed method, we analyze the household survey data on $n=4997$ adults between the ages of 16 and 64 in Mochudi, Botswana, out of whom 4045 (81%) had complete information on HIV testing. The majority of those who did not have HIV test results refused to participate in the HIV testing component. The baseline covariates include participant gender $(U_{1})$ and age in years $(U_{2})$, while the candidate outcome instrument $Z$ is a binary indicator of whether interviewer experience in years is in the top quartile. We implement the proposed estimator $\tilde{\phi}(c,d)=(\tilde{\mu},\tilde{\gamma})^{\mathrm{\scriptscriptstyle T}}$ with $c(y,x)=0$, $d(y,x)=yz$ based on the following main effects GLMs with canonical links, $\pi(y,x;\xi,\gamma)=\textup{expit}\{(1,z,u_{1},u_{2})\xi+\gamma y\}$, $p(Z=1\mid u;\beta)=\textup{expit}\{(1,u_{1},u_{2})\beta\}$, and $p(Y=1\mid u;\psi)=\textup{expit}\{(1,u_{1},u_{2})\psi\}$. These parametric models are chosen due to their simplicity for illustration, although in principle they can be checked against the observed data using goodness-of-fit tests.
For comparison, we also implement the estimator $\hat{\mu}_{{\textup{dr}}}$ of Sun et al., (2018) based on the GLM $p(Y=1\mid R=1,x;\psi_{1},\lambda)=\textup{expit}\{(1,u_{1},u_{2})\psi_{1}+\lambda z\}$, as well as the standard complete-case estimator $\hat{\mu}_{{\textup{cc}}}$ and the inverse propensity weighted estimator $\hat{\mu}_{{\textup{mar}}}$ based on propensity score estimated under the missing at random model, $\pi(y,x;\xi,\gamma=0)=\textup{expit}\{(1,z,u_{1},u_{2})\xi\}$. The analysis results are summarized in Table 2. The point estimates of HIV seroprevalence for $\tilde{\mu}$ and $\hat{\mu}_{{\textup{dr}}}$ are similar and substantially higher than those for $\hat{\mu}_{{\textup{cc}}}$ and $\hat{\mu}_{{\textup{mar}}}$, although at the expense of higher variance. This difference in efficiency reflects genuine uncertainty about the underlying nonresponse mechanism, because $\hat{\mu}_{{\textup{cc}}}$ and $\hat{\mu}_{{\textup{mar}}}$ impose a priori restrictions on the parameter space of $(\xi^{\mathrm{\scriptscriptstyle T}},\gamma)^{\mathrm{\scriptscriptstyle T}}$. The point estimate of $\tilde{\gamma}$ is $-1.854$ with 95% confidence interval $(-5.984,2.277)$, which suggests that HIV-infected persons are less likely to participate in the HIV testing component of the survey, although this difference is not statistically significant at the 0.05 significance level.
@@ 6 Extension to longitudinal studies with repeated outcome measures
We follow the longitudinal follow-up study design in Vansteelandt et al., (2007), in which the full data at each study cycle $1\leq t\leq T$ consists of $W_{t}=(Y_{t},{Z_{t},U_{t}})$, where $Y_{t}$ is an outcome, $Z_{t}$ an instrumental variable such as interviewer experience, and $U_{t}$ consists of other measured covariates. Suppose $X_{t}=(Z_{t},U_{t})$ is fully observed, but $Y_{t}$ is subject to missingness due to some subjects missing some study cycles. Let $R_{t}$ denote the occasion-specific missingness status. Thus, we only observe $O_{t}=(R_{t},R_{t}Y_{t},X_{t})$ at study cycle $t$. Here the nonresponse patterns can be nonmonotone, as $R_{t}=0$ does not necessarily imply that $R_{t+1}=0$. Let $\bar{O}_{t}=(O_{1},...,O_{t})$ denote the observed history up to and including cycle $t$, with $\bar{O}_{0}=\emptyset$. The nonresponse mechanism at each study cycle $t$ is captured by the occasion-specific extended propensity score $\pi_{t}(y,x,\bar{o}_{t-1})=p(R_{t}=1\mid Y_{t}=y,X_{t}=x,\bar{O}_{t-1}=\bar{o}_{t-1})$.
@@ 7 Discussion
In closing, we acknowledge certain limitations of the proposed semiparametric method and outline avenues for future work. The proposed doubly robust estimators can be improved in terms of efficiency (Tan,, 2006, 2010), bias (Vermeulen and Vansteelandt,, 2015) and robustness (Han and Wang,, 2013; Li et al.,, 2020). Our framework also opens the way to using various flexible nonparametric or machine learning methods to estimate the nuisance parameters (Chernozhukov et al.,, 2018, 2022). Although we have used the odds ratio, the proposed framework can in principle be extended to other measures of the residual association between the outcome of interest and nonresponse mechanism, for example on the multiplicative or additive scales. Lastly, nonignorable missing covariate data is also a long-standing problem in applied research, and methods for identification and inference abound. For example, Miao and Tchetgen Tchetgen, (2018) used shadow variables, while Bartlett et al., (2014) and Yang et al., (2019) assume that the covariate missingness mechanism is independent of the outcome. It will be of interest to explore the use of instrumental variables in similar settings with missing covariate data.
8 Acknowledgments

### v4

@@ 5 Illustration
To illustrate the proposed method, we analyze the household survey data on $n=4997$ adults between the ages of 16 and 64 in Mochudi, Botswana, out of whom 4045 (81%) had complete information on HIV testing. The majority of those who did not have HIV test results refused to participate in the HIV testing component. The baseline covariates include participant gender $(U_{1})$ and age in years $(U_{2})$, while the candidate outcome instrument $Z$ is a binary indicator of whether interviewer experience in years is in the top quartile. We implement the proposed estimator $\tilde{\phi}(c,d)=(\tilde{\mu},\tilde{\gamma})^{\mathrm{\scriptscriptstyle T}}$ with $c(w)=0$, $d(w)=yz$ based on the following main effects generalized linear models with canonical links, $\pi(w;\xi,\gamma)=\textup{expit}\{(1,z,u_{1},u_{2})\xi+\gamma y\}$, $p(Z=1\mid u;\beta)=\textup{expit}\{(1,u_{1},u_{2})\beta\}$, and $p(Y=1\mid u;\psi)=\textup{expit}\{(1,u_{1},u_{2})\psi\}$. These parametric models are chosen due to their simplicity for illustration, although in principle they can be checked against the observed data using goodness-of-fit tests.
For comparison, we also implement the estimator $\hat{\mu}_{{\textup{dr}}}$ of Sun et al. (2018) based on the working model $p(Y=1\mid R=1,x;\psi_{1},\lambda)=\textup{expit}\{(1,u_{1},u_{2})\psi_{1}+\lambda z\}$, as well as the standard complete-case estimator $\hat{\mu}_{{\textup{cc}}}$ and the inverse propensity weighted estimator $\hat{\mu}_{{\textup{mar}}}$ based on the missing at random propensity score model $\pi(w;\xi,\gamma=0)=\textup{expit}\{(1,z,u_{1},u_{2})\xi\}$. The analysis results are summarized in Table 2. The point estimates of HIV seroprevalence for $\tilde{\mu}$ and $\hat{\mu}_{{\textup{dr}}}$ are similar and substantially higher than those for $\hat{\mu}_{{\textup{cc}}}$ and $\hat{\mu}_{{\textup{mar}}}$, although at the expense of higher variance. This difference in efficiency reflects genuine uncertainty about the underlying nonresponse mechanism, because $\hat{\mu}_{{\textup{cc}}}$ and $\hat{\mu}_{{\textup{mar}}}$ impose a priori restrictions on the parameter space of $(\xi^{\mathrm{\scriptscriptstyle T}},\gamma)^{\mathrm{\scriptscriptstyle T}}$. The point estimate of $\tilde{\gamma}$ is $-1.854$ with 95% confidence interval $(-5.984,2.277)$, which suggests that HIV-infected persons are less likely to participate in the HIV testing component of the survey, although this difference is not statistically significant at the 0.05 significance level.
@@ 6 Extension to longitudinal studies with repeated outcome measures
We follow the longitudinal study design in Vansteelandt et al. (2007), in which the full data at each study cycle $1\leq t\leq T$ consists of $W_{t}=(Y_{t},{Z_{t},U_{t}})$, where $Y_{t}$ is an outcome, $Z_{t}$ an instrumental variable such as interviewer experience, and $U_{t}$ consists of other measured covariates. Suppose $X_{t}=(Z_{t},U_{t})$ is fully observed, but $Y_{t}$ is subject to missingness due to some subjects missing some study cycles. Let $R_{t}$ denote the occasion-specific missingness status. Thus, we only observe $O_{t}=(R_{t},R_{t}Y_{t},X_{t})$ at study cycle $t$. Here the nonresponse patterns can be nonmonotone, as $R_{t}=0$ does not necessarily imply that $R_{t+1}=0$. Let $\bar{O}_{t}=(O_{1},...,O_{t})$ denote the observed history up to and including cycle $t$, with $\bar{O}_{0}=\emptyset$. The nonresponse mechanism at each study cycle $t$ is captured by the occasion-specific extended propensity score $\pi_{t}(w_{t},\bar{o}_{t-1})=p(R_{t}=1\mid W_{t}=w_{t},\bar{O}_{t-1}=\bar{o}_{t-1})$.
@@ 7 Discussion
In closing, we acknowledge certain limitations of the proposed semiparametric method and outline avenues for future work. The proposed doubly robust estimators can be improved in terms of efficiency (Tan, 2006, 2010), bias (Vermeulen and Vansteelandt, 2015) and robustness (Han and Wang, 2013; Li et al., 2020). Our framework also opens the way to using various flexible nonparametric or machine learning methods to estimate the nuisance parameters (Chernozhukov et al., 2018, 2022). Although we have used the odds ratio, the proposed framework can in principle be extended to other measures of the residual association between the outcome of interest and nonresponse mechanism, for example on the multiplicative or additive scales. Lastly, nonignorable missing covariate data is also a long-standing problem in applied research, and methods for identification and inference abound. For example, Miao and Tchetgen Tchetgen (2018) used shadow variables, while Bartlett et al. (2014) and Yang et al. (2019) assume that the covariate missingness mechanism is independent of the outcome. It will be of interest to explore the use of instrumental variables in similar settings with missing covariate data.
Supplementary Materials
The online Supplementary Material contains proofs of propositions 1 and 2, a further discussion on local efficiency and additional simulation results.
Acknowledgements